# RAS: Retrieval-And-Structuring for Knowledge-Intensive LLM Generation

### v2

@@ Appendix A Limitations & Future Work
These limitations suggest several promising directions for future research, including improved knowledge extraction techniques and enhanced graph evolution strategies. A particularly promising direction is the integration of external knowledge graphs into RAS’s iterative loop, as its graph-based architecture naturally accommodates the incorporation of established KGs like Wikidata or domain-specific knowledge bases. This could significantly enhance the system’s reasoning capabilities by combining dynamically constructed knowledge with curated structured information. We summarize the potential future directions of extension in Figure 5.
Appendix B HotpotQA-SUBQ Dataset
The source of our dataset is HotpotQA [1].111We use the version hotpot_train_v1.1 from https://github.com/hotpotqa/hotpot. Our goal is to create a dataset that can let the model learn the capabilities of subquery generation and answering with intensive knowledge. We name the dataset as HotpotQA-SUBQ.
The construction of HotpotQA-SUBQ includes three steps: (1) filter out irrelevant/unnecessary “supporting docs” (§B.1), (2) generate one subquery for each supporting document (§B.2), and (3) sample labeling (§B.3).
B.1 Supporting Document Filtering
@@ B.4 Dataset Statistics
B.5 Data Examples
Appendix C Training Details
C.1 Text-to-Triples Model Training
For our text-to-triple conversion model, we fine-tuned a Flan-T5-Large model [69]222https://huggingface.co/pat-jj/text2triple-flan-t5 and a LLaMA-3.2-3B-Instruct model [70]333https://huggingface.co/pat-jj/text2graph-llama-3.2-3b to transform raw text passages into structured knowledge triples. The model processes input text sequences up to 512 tokens in length and generates structured triples in a standardized format “(S subject P predicate O object)”.
Our training dataset WikiOFGraph [44], curated by ChatGPT, comprises 5,851,776 text-triple pairs, with an additional 5,000 samples reserved for validation. Each training instance consists of a natural language text passage paired with its corresponding set of comma-separated triples. For example:
@@ C.1 Text-to-Triples Model Training
C.2 Multitask Training of Planning and Answering Capabilities
We employ Graph LLM [38, 11] as our foundational model architecture and utilize our constructed HotpotQA-SUBQ dataset to train a unified model capable of performing both graph-conditioned action planning and answer generation through multitask learning.
Since both tasks share an identical architectural foundation, as illustrated in Figure 1, we differentiate their functionality through specialized instruction sets. For Planning, we employ the following instruction template:
@@ C.2 Multitask Training of Planning and Answering Capabilities
For graph encoding, we implement Graph Transformer [65], selected for its robust capability in handling non-fully-connected graphs—a common characteristic of text-extracted triples. Our base language models comprise LLaMA-2-7B and LLaMA-3-8B, chosen both to maintain consistency with previous research [30, 31] and to investigate our framework’s performance scaling across different model capacities.
Our implementation of GraphLLM differs from G-Retriever [11] primarily due to the distinct nature of our graph structures. While G-Retriever operates on single interconnected graphs, our framework processes multiple potentially disconnected subgraphs, each corresponding to different subqueries. To address this architectural difference, we adopt a sequential encoding strategy: rather than encoding the entire graph at once, we process each subgraph individually using Graph Transformer, followed by mean pooling across all subgraph embeddings to produce the final encoded representation.
@@
Appendix D Evaluation Datasets & Metrics
D.1 Test Datasets
We evaluate RAS on diverse benchmark datasets spanning short-form QA, closed-set tasks, and long-form generation in the zero-shot setting, aligning with [30, 31]. Below we describe each dataset:
@@ D.1 Test Datasets
TriviaQA-unfiltered (TQA) [51]: A large-scale QA dataset containing 11,313 question-answer pairs in our test set. The questions are sourced from trivia enthusiasts and cover diverse topics.
2WikiMultiHopQA (2WQA) [53]: A multi-hop question answering dataset (with 12,576 samples in test set) that requires models to combine information from multiple Wikipedia articles to answer questions.
PopQA [52]: A dataset focusing on questions about long-tail entities, containing 1,399 queries where the monthly Wikipedia page views are less than 100. These questions test models’ ability to handle queries about less common entities.
@@
PubHealth (Pub) [54]: A dataset for verifying health-related claims. Models must classify statements as "true" or "false" based on scientific evidence. The dataset contains 987 test samples.
ARC-Challenge (ARC) [55]: A multiple-choice science question dataset designed to be challenging for models, requiring multi-hop reasoning and background knowledge. The dataset contains 1,172 test samples.
@@
ASQA [57]: A long-form question answering dataset that requires models to generate comprehensive answers with proper citation of sources. Models must balance information completeness with factual accuracy. The dataset contains 948 test samples.
ELI5 [58]: A dataset derived from the "Explain Like I’m Five" subreddit, containing questions seeking straightforward explanations of complex topics. The dataset contains 1,000 test samples.
D.2 Evaluation Metrics
@@
For PopQA and TriviaQA, we use the "golden match" metric [30, 71, 8]444Implementation aligned with https://github.com/AkariAsai/self-rag/blob/main/retrieval_lm/metrics.py, where a prediction is considered correct if it contains any normalized version of the ground truth answers :
@@
For 2WikiMultiHopQA, we follow RPG [31]555Implementation aligned with https://github.com/haruhi-sudo/RPG/blob/main/retriever/src/evaluation.py to use token-level F1 score between prediction and ground truth :
@@
Closed-set Tasks: For both PubHealth and ARC-Challenge, we use accuracy [30], computed as:
@@
Long-form Generation: For ASQA and ELI5, we use multiple metrics666Implementation aligned with https://github.com/princeton-nlp/ALCE/blob/main/eval.py:
@@
Appendix E Inference & Evaluation Details
@@
For the open-source models, we first fine-tuned the GraphLLM architecture on our HotpotQA-SUBQ dataset (see Section C.2). We then conducted zero-shot knowledge-intensive generation tests. With Claude-3.5-Sonnet, we used few-shot prompting for both action planning and answer generation phases. Below, we detail our approach for each model type.
E.1 Closed-Source Model Settings
@@
In addition, for ASQA and ELI5 with closed-source models (RAS and all the other baselines (e.g., SuRe) using Sonnet-3.5), we conduct few-shot inference with two in-context learning demonstrations, aligning with previous study’s [56] implementation.777https://github.com/princeton-nlp/ALCE/blob/main/run.py888Exemplars can be found at https://github.com/princeton-nlp/ALCE/tree/main/prompts
E.2 Open-Source Model Settings
For inference with open-source model, we use our GraphLLM trained by Hotpot-SUBQ (see Appendix C.2), we use 8 NVIDIA RTX 6000 Ada Generation with 48GB memory and CUDA version 12.4. We use Python 3.10, PyTorch 2.1.2, and torch-geometric 2.6.1 throughout all experiments.
For more efficient text retrieval, we split the dense index of the corpus into eight splits, and load them on eight GPUs with faiss-gpu 1.7.2.999All the Wikipedia dumps (2018, 2020) we used are downloaded from https://github.com/facebookresearch/atlas
We set maximum new tokens as 100 for PopQA, TriviaQA, and 2WikiMultihopQA, as 50 for PubHealth and ARC-Challenge, and as 300 for ASQA and ELI5, aligning with previous study [30, 31]. All the generation are configured with batch size of 1.
As we always set the main query as the first subquery, we pre-extract the triples from the context at the first iteration using our text-to-triple model for batch triple extraction.
E.3 Instructions
@@
Appendix F Supplementary Results and Analysis
@@
We present two representative cases where the quality of retrieved structured knowledge significantly impacts final answer generation. These examples are distinct from the qualitative examples shown in Section G and highlight both successful and failed reasoning paths.
@@
Appendix G Qualitative Examples
@@
Appendix H Hyperparameter Study
@@
| Max new tokens | Task-specific (see Appendix E.2) |
@@
Appendix I Broader Impacts, Safeguards, and Assets
Broader Impacts. The RAS framework presents several promising impacts for both research and practical applications. By transforming unstructured retrieved content into structured, query-specific knowledge graphs, RAS improves reasoning transparency and factual reliability—qualities that are especially valuable in education, scientific research, and technical writing. The ability to explicitly organize retrieved knowledge into interpretable structures also opens up new possibilities for human-in-the-loop AI systems, enabling users to trace and verify model reasoning paths.

### v3

@@ Appendix A Limitations & Future Work
These limitations suggest several promising directions for future research, including improved knowledge extraction techniques and enhanced graph evolution strategies. A particularly promising direction is the integration of external knowledge graphs into RAS’s iterative loop, as its graph-based architecture naturally accommodates the incorporation of established KGs like Wikidata or domain-specific knowledge bases. This could significantly enhance the system’s reasoning capabilities by combining dynamically constructed knowledge with curated structured information.
Appendix B The Use of Large Language Models
In this work, large language models (LLMs) were used solely for drafting and polishing the manuscript. Their role was limited to grammar correction, phrasing refinement, and enhancing readability. All research ideas, methodological innovations, theoretical analyses, and experimental work were entirely conceived, designed, and carried out by the authors. The authors take full responsibility for the scientific content of the paper.
Appendix C HotpotQA-SUBQ Dataset
The source of our dataset is HotpotQA (Yang et al., 2018).111We use the version hotpot_train_v1.1 from https://github.com/hotpotqa/hotpot. Our goal is to create a dataset that can let the model learn the capabilities of subquery generation and answering with intensive knowledge. We name the dataset as HotpotQA-SUBQ.
The construction of HotpotQA-SUBQ includes three steps: (1) filter out irrelevant/unnecessary “supporting docs” (§C.1), (2) generate one subquery for each supporting document (§C.2), and (3) sample labeling (§C.3).
C.1 Supporting Document Filtering
@@ C.4 Dataset Statistics
C.5 Data Examples
Appendix D Training Details
D.1 Text-to-Triples Model Training
For our text-to-triple conversion model, we fine-tuned a Flan-T5-Large model (Chung et al., 2024) and a LLaMA-3.2-3B-Instruct model (Grattafiori et al., 2024) to transform raw text passages into structured knowledge triples. The model processes input text sequences up to 512 tokens in length and generates structured triples in a standardized format “(S subject P predicate O object)”.
Our training dataset WikiOFGraph (Kim et al., 2024a), curated by ChatGPT, comprises 5,851,776 text-triple pairs, with an additional 5,000 samples reserved for validation. Each training instance consists of a natural language text passage paired with its corresponding set of comma-separated triples. For example:
@@ D.1 Text-to-Triples Model Training
D.2 Multitask Training of Planning and Answering Capabilities
We employ Graph LLM (Perozzi et al., 2024; He et al., 2024) as our foundational model architecture and utilize our constructed HotpotQA-SUBQ dataset to train a unified model capable of performing both graph-conditioned action planning and answer generation through multitask learning.
Since both tasks share an identical architectural foundation, as illustrated in Figure LABEL:fig:overviewLABEL:fig:overview, we differentiate their functionality through specialized instruction sets. For Planning, we employ the following instruction template:
@@ D.2 Multitask Training of Planning and Answering Capabilities
For graph encoding, we implement Graph Transformer (Shi et al., 2020), selected for its robust capability in handling non-fully-connected graphs—a common characteristic of text-extracted triples. Our base language models comprise LLaMA-2-7B and LLaMA-3-8B, chosen both to maintain consistency with previous research (Asai et al., 2023; Lyu et al., 2024b) and to investigate our framework’s performance scaling across different model capacities.
Our implementation of GraphLLM differs from G-Retriever (He et al., 2024) primarily due to the distinct nature of our graph structures. While G-Retriever operates on single interconnected graphs, our framework processes multiple potentially disconnected subgraphs, each corresponding to different subqueries. To address this architectural difference, we adopt a sequential encoding strategy: rather than encoding the entire graph at once, we process each subgraph individually using Graph Transformer, followed by mean pooling across all subgraph embeddings to produce the final encoded representation.
@@
Appendix E Evaluation Datasets & Metrics
E.1 Test Datasets
We evaluate RAS on diverse benchmark datasets spanning short-form QA, closed-set tasks, and long-form generation in the zero-shot setting, aligning with (Asai et al., 2023; Lyu et al., 2024b). Below we describe each dataset:
@@ E.1 Test Datasets
TriviaQA-unfiltered (TQA) (Joshi et al., 2017): A large-scale QA dataset containing 11,313 question-answer pairs in our test set. The questions are sourced from trivia enthusiasts and cover diverse topics.
2WikiMultiHopQA (2WQA) (Ho et al., 2020): A multi-hop question answering dataset (with 12,576 samples in test set) that requires models to combine information from multiple Wikipedia articles to answer questions.
PopQA (Mallen et al., 2022): A dataset focusing on questions about long-tail entities, containing 1,399 queries where the monthly Wikipedia page views are less than 100. These questions test models’ ability to handle queries about less common entities.
@@
PubHealth (Pub) (Zhang et al., 2023a): A dataset for verifying health-related claims. Models must classify statements as "true" or "false" based on scientific evidence. The dataset contains 987 test samples.
ARC-Challenge (ARC) (Clark et al., 2018): A multiple-choice science question dataset designed to be challenging for models, requiring multi-hop reasoning and background knowledge. The dataset contains 1,172 test samples.
@@
ASQA (Stelmakh et al., 2022): A long-form question answering dataset that requires models to generate comprehensive answers with proper citation of sources. Models must balance information completeness with factual accuracy. The dataset contains 948 test samples.
ELI5 (Fan et al., 2019): A dataset derived from the "Explain Like I’m Five" subreddit, containing questions seeking straightforward explanations of complex topics. The dataset contains 1,000 test samples.
E.2 Evaluation Metrics
@@
For PopQA and TriviaQA, we use the "golden match" metric (Asai et al., 2023; Min et al., 2019; Guu et al., 2020)222Implementation aligned with https://github.com/AkariAsai/self-rag/blob/main/retrieval_lm/metrics.py, where a prediction is considered correct if it contains any normalized version of the ground truth answers :
@@
For 2WikiMultiHopQA, we follow RPG (Lyu et al., 2024b)333Implementation aligned with https://github.com/haruhi-sudo/RPG/blob/main/retriever/src/evaluation.py to use token-level F1 score between prediction and ground truth :
@@
Closed-set Tasks: For both PubHealth and ARC-Challenge, we use accuracy (Asai et al., 2023), computed as:
@@
Long-form Generation: For ASQA and ELI5, we use multiple metrics444Implementation aligned with https://github.com/princeton-nlp/ALCE/blob/main/eval.py:
@@
Appendix F Inference & Evaluation Details
@@
For the open-source models, we first fine-tuned the GraphLLM architecture on our HotpotQA-SUBQ dataset (see Section D.2). We then conducted zero-shot knowledge-intensive generation tests. With Claude-3.5-Sonnet, we used few-shot prompting for both action planning and answer generation phases. Below, we detail our approach for each model type.
F.1 Closed-Source Model Settings
@@
In addition, for ASQA and ELI5 with closed-source models (RAS and all the other baselines (e.g., SuRe) using Sonnet-3.5), we conduct few-shot inference with two in-context learning demonstrations, aligning with previous study’s (Gao et al., 2023) implementation.555https://github.com/princeton-nlp/ALCE/blob/main/run.py666Exemplars can be found at https://github.com/princeton-nlp/ALCE/tree/main/prompts
F.2 Open-Source Model Settings
For inference with open-source model, we use our GraphLLM trained by Hotpot-SUBQ (see Appendix D.2), we use 8 NVIDIA RTX 6000 Ada Generation with 48GB memory and CUDA version 12.4. We use Python 3.10, PyTorch 2.1.2, and torch-geometric 2.6.1 throughout all experiments.
For more efficient text retrieval, we split the dense index of the corpus into eight splits, and load them on eight GPUs with faiss-gpu 1.7.2.777All the Wikipedia dumps (2018, 2020) we used are downloaded from https://github.com/facebookresearch/atlas
We set maximum new tokens as 100 for PopQA, TriviaQA, and 2WikiMultihopQA, as 50 for PubHealth and ARC-Challenge, and as 300 for ASQA and ELI5, aligning with previous study (Asai et al., 2023; Lyu et al., 2024b). All the generation are configured with batch size of 1.
As we always set the main question as the first subquery, we pre-extract the triples from the context at the first iteration using our text-to-triple model for batch triple extraction.
F.3 Instructions
@@
Appendix G Cost & Efficiency Comparison with Other Graph-based RAG Methods
Static vs. Dynamic Structuring.
Static corpus-level graph methods such as GraphRAG (Edge et al., 2024) and HippoRAG (Gutiérrez et al., 2024) require building and maintaining large knowledge graphs offline. This involves millions of LLM calls for entity/relation extraction and summarization, plus heavy clustering and embedding computation. At the scale of Wikipedia 2018 (roughly 5M articles, 3B tokens), these pipelines incur substantial costs, days of GPU/CPU compute, and 50 GB storage overhead. In contrast, our RAS constructs question-specific knowledge graphs dynamically at inference, eliminating expensive offline indexing. The only extra step is lightweight triple extraction from retrieved passages, which we implement with a small LLaMA-3B model accelerated by vLLM (Kwon et al., 2023), achieving throughput of nearly 5K tokens/s (Table 3).
Cost Estimation Methodology.
We estimate costs based on empirical data from Microsoft’s GraphRAG analysis 888https://techcommunity.microsoft.com/blog/azure-ai-foundry-blog/graphrag-costs-explained-what-you-need-to-know/4207978 and scale to Wikipedia 2018. Cost estimates assume current API pricing and may vary significantly based on implementation details, model selection, and optimization strategies.
Computation of Indexing Cost.
We estimate costs using the following assumptions:
Corpus size: Wikipedia 2018 3B tokens, or 5M articles with 600 tokens/article.
GraphRAG pipeline: Based on Microsoft’s empirical analysis showing $0.000011 per word for GPT-4o-mini.
For 3B tokens (2.25B words), base processing cost: $25k–$30k
Entity/relation extraction with few-shot prompting increases token usage by 3–5×: $75k–$150k
Community summarization for 10k communities at $0.02 per community: $200–$500
Additional overhead (clustering, embeddings, storage): $10k–$25k
Total GraphRAG: $85k–$175k for GPT-4o-mini; $250k–$500k for GPT-4o
HippoRAG pipeline: More efficient extraction but still corpus-wide processing.
Entity extraction: 3B tokens at reduced complexity $15k–$30k
Synonym linking and embedding overhead: $5k–$15k
Graph construction and indexing: $5k–$10k
Total HippoRAG: $25k–$55k
RAS pipeline: No offline corpus-wide indexing. Triple extraction uses local LLaMA-3B model.
Offline indexing cost: $0 (no corpus preprocessing required)
Runtime triple extraction: Local model inference (negligible marginal cost)
Total RAS: $0 offline costs
Cost Sensitivity Analysis.
Costs vary significantly based on:
Model selection: GPT-4o costs 3× more than GPT-4o-mini
Chunk size: Smaller chunks (300 vs 1200 tokens) can reduce costs by 30–50%
Prompt optimization: Well-tuned prompts can reduce token usage by 20–40%
Processing parameters: Conservative estimates assume non-optimized settings
Quantitative Comparison.
Table 6 summarizes estimated costs with uncertainty ranges.
| Method | Offline LLM Cost | Compute Time | Storage | Ready for RAG? |
| GraphRAG | $85k–$500k (model-dependent) | >7 days (distributed) | 50–70 GB | Yes (community reports) |
| HippoRAG | $25k–$55k (optimized extraction) | >3 days (cluster) | 30–50 GB | Yes (node-passage matrix) |
| RAS (ours) | Negligible | Inference-time only | Small (triple cache) | Yes (dynamic graph) |
Efficiency Visualization.
We visualize the contrast in offline indexing cost ranges. RAS requires no precomputation, while GraphRAG and HippoRAG incur substantial upfront costs.
Takeaway.
RAS achieves the structured reasoning benefits of graphs without any offline indexing costs, compared to GraphRAG ($85k–$500k) and HippoRAG ($25k–$55k). By eliminating upfront investment requirements and enabling pay-per-query scaling through local model inference, RAS makes graph-enhanced RAG accessible for scenarios where large-scale preprocessing is impractical or cost-prohibitive.
Appendix H Supplementary Results and Analysis
@@
We present two representative cases where the quality of retrieved structured knowledge significantly impacts final answer generation. These examples are distinct from the qualitative examples shown in Section I and highlight both successful and failed reasoning paths.
@@
Appendix I Qualitative Examples
@@
Appendix J Hyperparameter Study
@@
| Max new tokens | Task-specific (see Appendix F.2) |
@@
Appendix K Broader Impacts, Safeguards, and Assets
Broader Impacts. The RAS framework presents several promising impacts for both research and practical applications. By transforming unstructured retrieved content into structured, question-specific knowledge graphs, RAS improves reasoning transparency and factual reliability—qualities that are especially valuable in education, scientific research, and technical writing. The ability to explicitly organize retrieved knowledge into interpretable structures also opens up new possibilities for human-in-the-loop AI systems, enabling users to trace and verify model reasoning paths.